Activity-Based Costing Slack Corporation has the following predicted indirect costs and cost drivers for 2016 for the given activity cost pools Fabrication Department Finishing Department Cost Driver Maintenance Materials handling Machine setups Inspections $35,000 30,000 70,000 $25,000 Machine hours 15,000 Material moves 5,000 Machine setups 25,000 Inspection hours $135,000 $70,000 The following activity predictions were also made for the year: Machine hours Material moves Machine setups Inspection hours Fabrication Department Finishing Department 5,000 1,500 50 1,000 10,000 3,000 700 It is assumed that the cost per unit of activity for a given activity does not vary between departments Slacks president, Charles Slack, is trying to evaluate the companys product mix strategy regarding two of its five product models, Zx300 and SL500. The company has been using a company-wide overhead rate based on machine hours but is considering switching to either department rates or activity-based rates The production manager has provided the following data for the production of a batch of 100 units for each of these models: ZX300 SL500 Direct materials cost Direct labor cost Machine hours (Fabrication) Machine hours (Finishing) Materials moves Machine setups Inspection hours $14,000 $20,000 $5,000 $4,000 700 100 50 500 200 30 30 60a. Determine the cost of one unit each of ZX300 and SL500, assuming a company-wide overhead rate is used based on total machine hours. Round rate to two decimal places. Overhead rate based on total machine hours 0 Use rounded overhead rate calculated above for calculations below. Round cost answers to the nearest whole number. Round cost per unit to two decimal places. per machine hour Product Costs per Unit ZX300 SL500 Direct materials 0 Direct labor Manufacturing overhead: 0 Total cost per batch Number of units per batch 0 Cost per unit b. Determine the cost of one unit of ZX300 and SL500, assuming department overhead rates are used. Overhead is assigned based on machine hours in both departments. Round rate to two decimal places, if applicable Overhead rate based on machine hours for Fabrication Dept. $0 per machine hour Overhead rate based on machine hours for Finishing Dept. $ 0 per machine hour Use rounded overhead rate calculated above for calculations below. Round cost answers to the nearest whole number, when needed. Round cost per unit to two decimal places, if needed Product Costs per Unit Product Costs per Unit SL500 Direct materials 0 Direct labor Manufacturing overhead: Fabrication Dept. Finishing Dept. Total cost per batch 0 Number of units per batch Cost per unitc. Determine the cost of one unit of ZX300 and SL500, assuming activity-based overhead rates are used for maintenance, materials handling, machine setup, and inspection activities Round rate to two decimal places, if applicable Activity-based overhead rates: Maintenance 0 per machine hour 0 per materials move 0 per machine setup 0 per inspection hour Materials handling Machine setup Inspection activities Use rounded overhead rate calculated above for calculations below. Round cost answers to the nearest whole number, when needed. Round cost per unit to two decimal places, if needed Product Costs per Unit ZX300 SL500 Direct materials 0 Direct labor Manufacturing overhead Maintenance activity Materials handling activity Machine setups activity 0 0 0 0 Inspections activity 0 Total cost per batch 0 Number of units per batch 0 0 Cost per unit
Answer #1

a)Overhead rate=Total manufacturing overhead/total machine hours
=(135000+70000)/(10000+5000)
=13.67

b) overhead rate based on machinf hrs for fabrication dept=department overhead/department machine hours
=135000/10000=13.5
finishing department=70000/5000=14

c)Activity based rates:
Maintenace=total costs/total machine hours
=(35000+25000)/(10000+5000)=4
Materials handling=(30000+15000)/(3000+1500)=10
Machine setup=(70000+5000)/(700+50)=100
Inspection activites=25000/1000=25

Overhead rate 13.6666667 Zx300 (135000+70000)/(10000+ Zx300 14000 5000 (500+200) B1 SUM(B3:B5) 1 Overhead rate SL500 SL500 Di